# AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL

### v2

@@
Abstract
Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime. Although both approaches have shown promise, they also face critical limitations. LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient. On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience. In this paper, we present AQORA, a learned adaptive query optimizer that continuously refines query plans through stage-level feedback, unifying pre-execution and in-execution optimization within a single learning process. AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration. Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
I Introduction
Query optimization plays a central role in database systems, determining the physical plan for a given query. Typically, this process takes place prior to query execution. The query optimizer selects a fixed execution plan based on estimated data statistics, analytical cost models, and a predefined set of transformation rules. Once execution begins, the query processing engine adheres strictly to the chosen plan, without the opportunity to revise it based on actual runtime behavior.
In response to the limitations of static, pre-execution query planning, subsequent research has evolved along two primary directions. One line of work focuses on developing learning-based algorithms that assist or replace the traditional query optimizer in making key planning decisions (such as join ordering, operator selection, or cardinality estimation) prior to execution beginning[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. These learned query optimizers (LQOs) leverage historical execution data to enhance decision-making and generalize to unseen queries. In parallel, another line of research shifts part of the optimization responsibility into the execution phase, enabling dynamic plan adaptation based on actual runtime statistics[13, 14, 15, 16, 17, 18, 19, 20]. This paradigm, known as adaptive query processing (AQP), has been increasingly adopted by modern engines such as Spark SQL[21, 22], allowing certain decisions to be deferred until runtime.
While both approaches have significantly advanced the state of query optimization, they each face inherent limitations stemming from their core methodology. For LQOs, their optimization is confined to the pre-execution phase, which gives rise to two fundamental challenges:
C1: Lack of runtime information. Pre-execution query optimization requires the system to make planning decisions without access to essential runtime information, such as the actual cardinalities of intermediate results. Although researchers have proposed various encoding methods and predictive models to improve pre-execution estimation, these approaches still often fail to consistently outperform traditional rule-based optimizers such as PostgreSQL’s planner [23].
C2: Lack of fine-grained feedback. Most LQOs receive only a single final feedback metric (e.g., end-to-end execution time) for each query, even though they make multiple interdependent optimization decisions before execution. This coarse-grained signal fails to identify which specific decision led to performance changes. More critically, LQOs cannot observe intermediate execution states and revise their decisions based on actual runtime behavior.
On the other hand, prior work [24] has shown that even simple AQP techniques can, in some cases, match or even outperform more sophisticated LQOs. However, AQP methods also face two fundamental challenges:
C3: Lack of flexible adaptation. Most existing AQP methods rely on static, rule-based heuristics to trigger runtime adjustments, such as switching join strategies or enabling dynamic repartitioning. These rules are often tuned manually via parameters set by DBAs, which limits their flexibility and generalization to unseen workloads. Moreover, such methods rarely consider join reordering, despite it being one of the most impactful aspects of query optimization.
C4: Lack of practical deployability. Although AQP methods offer many valuable ideas, only a small number have been implemented in real-world database systems. A main barrier is the inherent complexity of combining planning and execution, which complicates testing, debugging, and performance tuning. Introducing learning-based mechanisms in this context further increases system complexity, making reliable deployment even more challenging.
This leads to a fundamental question: can we design a query optimizer that continuously learns and adapts across both the pre- and in-execution phases? Fig. 1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs. in-execution) and how decisions are made (rule-based vs. learning-based). Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback. In contrast, AQP methods, such as Spark SQL’s adaptive query execution and runtime filters, perform limited rule-based adjustments during execution. As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned adaptive query optimizer that performs in-execution plan refinement on Spark SQL. It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision. It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
Figure 1: AQORA is the first learned query optimizer achieving execution-time query plan optimization on Spark SQL.
AQORA addresses the aforementioned challenges through four key strategies:
S1: Runtime plan feature representation. AQORA removes the constraint of making all optimization decisions before execution begins. Instead, it dynamically collects runtime signals, such as actual cardinalities and execution statistics, and uses them as query plan features. This enables the model to make more informed and adaptive decisions based on real-time execution context. Compared to approaches that rely solely on pre-execution estimation [25, 26], this design also reduces the reliance on handcrafted features and complex estimators while improving robustness and accuracy.
@@ I Introduction
S3: Learning-based adaptive strategies formulation. Instead of relying on fixed, rule-based heuristics, AQORA’s decision model automatically learns its optimization strategies from feedback collected during query execution. This data-driven approach enhances generalization across diverse query workloads without requiring manual tuning. To enhance learning stability and final performance, AQORA adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner. In addition, AQORA supports a two-phase optimization mechanism, allowing strategies learned during execution to be reused in the planning phase of future queries.
S4: Lightweight and seamless system integration. Unlike prior approaches [10, 27] that require intrusive source code modifications, AQORA adopts a non-intrusive, plugin-style architecture. It can be enabled through simple configuration adjustments without altering the underlying query engine’s codebase. This plug-and-play design significantly reduces integration overhead and enables seamless deployment through minimal configuration.
Our contributions are summarized as follows:
  • We propose AQORA, an adaptive query optimizer for Spark SQL that integrates reinforcement learning into Spark’s adaptive query execution framework. AQORA comprises a decision model based on actor-critic learning and a planner extension that enables online plan modification during execution.
@@
  • We adopt a lightweight decision model built upon the TreeCNN architecture [28], which learns optimization actions directly from stage-level feedback collected during query execution. By using true cardinalities as input features (and avoiding sample bitmaps or histogram encodings), the model achieves better query performance with a simplified architecture.
@@
  • We design a plug-and-play planner extension built upon the extensibility interfaces of modern query engines, enabling non-intrusive integration with existing deployments. The extension intercepts runtime execution plans, interacts with the decision model, and applies optimization actions such as join reordering.
@@
  • We conduct extensive experiments on three public benchmarks. Results show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
II Related Work
In this section, we review related work on LQO and AQP. Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
Plan-constructor
@@ II Related Work
Plan-constructor methods aim to build query plans directly from scratch. These approaches enable the discovery of novel and potentially more efficient plans that lie beyond the capabilities of traditional optimizers. Neo [1] pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle. Balsa [2] further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques. LOGER [3] enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly. GLO [4] improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters. However, constructing plans from scratch suffers from two fundamental limitations: 1) the initial search often begins with poorly performing plans; and 2) the search space is extremely large, making efficient convergence challenging. In contrast, AQORA addresses these issues by starting from a reasonable initial plan generated by the native optimizer and learning to make incremental adjustments within a constrained action space.
Plan-steerer
@@
Plan-steerer methods improve query plans by exploiting the internal mechanisms of the native optimizer. Instead of constructing plans from scratch, they guide the optimizer to explore multiple candidate plans and select the one with the lowest estimated cost. For example, Bao [5], DeepO [8], SCOPE [6], FASTgres [7], and AutoSteer [9] generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers. HybridQO [11] explores alternative join orders by injecting leading hints. Lero[10] modifies cardinality estimates in the native optimizer to produce diverse plans. LEON [12] enumerates sub-plans directly using the native optimizer, while FOSS [30] refines plans after initial optimization by fixing join orders and methods. However, these methods rely heavily on the internal mechanisms of specific DBMS-native optimizers. As a result, the effectiveness of some of these methods is limited when applied to different DBMSes.
@@
AQP has been extensively studied to enable query plan adjustments during execution based on runtime feedback [13, 14, 15, 16, 18, 17, 19, 20, 27]. For example, Lookahead Information Passing [19] generalizes semi-join techniques to enhance the pipelining of equijoins. Adaptive Join Algorithm [31] dynamically switches between hash join and nested loop join during runtime. UDAO [27] adjusts parameters at multiple levels of granularity throughout query execution. RPT [32] extends predicate transfer with efficient semi-join reduction and join order validation to achieve provably robust join ordering for acyclic queries. In contrast to these methods, AQORA introduces the first learning-based AQP framework capable of continuously improving from experience and adaptively optimizing joins at runtime.
III Preliminary
In this section, we formulate the problem of query optimization and provide background on Spark SQL’s current approach.
III-A The Problem of Query Optimization
Query optimization involves many considerations, including predicate pushdown, expression simplification, and join optimization, etc. In this paper we focus on two core decisions: (1) determining the optimal join order and (2) selecting appropriate physical operators.
Consider a query $Q$ that joins $n$ tables $t_{1},t_{2},\cdots,t_{n}$}. If we restrict plan space to left-deep trees (Fig. 2(a)), there are up to $n!$ possible join orders. Allowing bushy trees (Fig. 2b) enlarges the search space to (2n2)!/(n1)!(2n-2)!/(n-1)! [33].
(a) left-deep join plan
(b) bushy join plan
Figure 2: In a left-deep tree plan, the left child of any join operation is a subtree, and the right child must be a leaf node. In contrast, bushy tree plans allow both children of any join operation to be subtrees
Join order and operator selection are critical factors in determining query execution efficiency. The core challenge of query optimization is to find a query plan $p$ that specifies both a join order and a set of physical operators, while minimizing the overall execution cost, which can be expressed as follows:
$\arg\min_{p}C(p)=C_{\text{plan}}(p)+C_{\text{execute}}(p)$
Here, $C_{\text{plan}}$ denotes the time spent generating candidate plans and identifying the final plan $p$, while $C_{\text{execute}}$ represents the actual execution time of plan $p$.
III-B Cost Based Optimization
Spark SQL’s cost-based optimizer (CBO) selects join orders using a dynamic programming algorithm that relies on cardinality estimates. This approach performs reasonably well for simple queries, where even with inaccurate cardinality estimates, suboptimal join orders still keep the planning cost ($C_{\text{plan}}$) within acceptable bounds. However, it encounters significant challenges when handling complex queries with many joins. The exponential time complexity of dynamic programming leads to rapid growth in optimization time, often resulting in both prolonged $C_{\text{plan}}$ and suboptimal plan selection. As shown in Fig. 3, $C_{\text{plan}}$ can become the dominant component of total cost as the number of joins increases. In fact, for JOB query 29a, the execution time shows little difference between enabling and disabling CBO, whereas the optimization time dominates when CBO is enabled. While we acknowledge that this issue may be specific to Spark SQL’s CBO implementation, our goal is not to replace or circumvent CBO but to provide a practical mechanism that dynamically mitigates such performance pitfalls.
Figure 3: Execution and optimization times of selected JOB queries with and without CBO.
CBO determines which join operator to use during physical optimization based on cardinality estimates. A key configuration is autoBroadcastJoinThreshold (BJT), which controls when broadcast joins are applied. If either side of the join is estimated to be smaller than BJT, Spark uses a broadcast hash join; otherwise, it falls back to shuffle joins such as sorted merge join.
However, this approach can lead to two problems. First, inaccurate cardinality estimates may result in selecting an inappropriate join type. Second, if the actual data size exceeds BJT but the estimate is incorrect, Spark may attempt to broadcast a large table, potentially causing worker crashes due to memory overload. To prevent such crashes, administrators often set BJT to a conservative value. While this safety measure enhances stability, it also limits the performance benefits offered by broadcast joins.
III-C Adaptive Query Execution
Spark SQL 3.0 (2020) introduced adaptive query execution (AQE), which is an implementation of AQP and aims to tackle the challenges of cardinality estimation. As Xue et al. [22] claimed, maintaining accurate statistics in distributed systems like Spark is more difficult than in traditional databases because the continuous flow of data makes real-time statistics maintenance impractical. However, abandoning cardinality estimation is not feasible, as optimization without statistical guidance tends to perform even worse. AQE’s solution is to defer certain decisions (like determine join operators): it starts with rough estimates and then re-optimizes the query during execution using actual runtime metrics such as shuffle partition sizes, as illustrated in Fig. 4.
Figure 4: This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
AQE leverages runtime statistics to improve broadcast join decisions and adjust physical operators. It also performs specific optimizations such as merging shuffle partitions and fixing skewed joins. However, AQE cannot modify the initial join order determined during logical optimization and lacks the capability to learn from past executions.
IV System Overview
In this section, we introduce the main components of AQORA and its basic workflow.
As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The figure illustrates an example where AQORA performs an optimization step during the execution of the STACK query q1#1.

### v3

# LQRS: Learned Query Re-optimization Framework for Spark SQL

@@
Abstract.
The query optimizer is a fundamental component of database management systems. Recent studies have shown that learned query optimizers outperform traditional cost-based query optimizers. However, they fail to exploit valuable runtime observations generated during query execution to dynamically re-optimize the plan, thereby limiting further improvements in query performance. To address this issue, we propose learned query re-optimization, which allows optimization decisions to be deferred to execution time and guided by actual runtime observations. We realize this idea through LQRS, a learned query re-optimization framework that builds upon Spark SQL, exploiting runtime observations for dynamic plan refinement. Specifically, LQRS employs a curriculum reinforcement learning strategy and jointly supports pre-execution and in-execution optimization, allowing knowledge learned during execution to directly benefit pre-execution planning. Furthermore, we design a plug-and-play planner extension built upon the extensibility interfaces of Spark SQL, enabling online plan modification. Experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/heyewuyue1/LQRS.
1. Introduction
The query optimizer is a fundamental component of modern database management systems. Its primary role is to select an efficient execution plan for each query. Traditional cost-based query optimization (Selinger et al., 1979) is a common technique in query optimizers, where the idea is to select the plan with the minimum estimated cost. However, these cost models employ a variety of handcrafted formulas to approximate actual execution costs; such approximations are often inaccurate, resulting in suboptimal execution plans.
To address this problem, people propose to employ machine learning in query optimization (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021; Chen et al., 2024b; Zhong et al., 2024; Sun et al., 2022; Yu et al., 2022; Chen et al., 2023b; Marcus et al., 2019; Yang et al., 2022; Chen et al., 2023a, 2024a). Such a learned query optimizer (LQO) leverages models trained on historical execution feedback to improve the enumeration and evaluation of candidate execution plans, thereby enabling more effective optimization decisions.
@@ 1. Introduction
Table 1. STACK Q1#1 involves optimizing a join plan over $t$, $s$, $q$, and tqtq. Lero fixes the join order before execution, whereas LQRS adjusts the join order based on runtime information, resulting in a shorter execution time.
Plan Lero (Chen et al., 2024b) LQRS (ours)
Pre-execution $(t\bowtie tq)\bowtie(s\bowtie q)$ $q\bowtie tq\bowtie t\bowtie s$
In-execution Fixed during execution $t\bowtie tq\bowtie q\bowtie s*$
Execution time 21.52s 15.24s
    During execution, LQRS observes that the scan on table $t$ produces only one tuple and reorders the join accordingly.
Such a limitation goes beyond LQO’s expressiveness or feature engineering. For example, characteristics such as the value distribution of aggregation keys or the exact selectivity of complex predicates only manifest once data is actually processed (Hellerstein et al., 1997). Even with sophisticated encodings and learned estimators, LQO models restricted to pre-execution optimization frequently struggle to produce reliable predictions for such properties, and empirical studies have shown that learned estimators often fail to consistently outperform traditional cost-based optimizers such as PostgreSQL’s planner (Heinrich et al., 2025). This fact suggests that the challenge arises not merely from modeling limitations, but from making decisions without access to critical runtime observations. This motivated us to think about an important question: how can we design a learned query re-optimization framework that can collect runtime observations and use them to dynamically re-optimize previously generated query plans during query execution?
To address this issue, we propose learned query re-optimization, which allows optimization decisions to be deferred to execution time and guided by actual runtime observations. We realize this idea through LQRS, a learned query re-optimization framework that exploits runtime observations for dynamic plan refinement. Our realization of LQRS builds upon the Adaptive Query Execution (AQE) foundation of Spark SQL (Xue et al., 2024), which provides the necessary runtime hooks for collecting execution feedback and revising optimization decisions. As illustrated in Figure 1, existing query optimization techniques intervene at different phases of a query’s lifecycle: traditional rule-based operator optimization and cost-based join reordering are performed entirely before execution; most learned query optimizers such as Lero(Chen et al., 2024b) and AutoSteer (Anneser et al., 2023) are also restricted to the pre-execution phase; while adaptive query processing techniques like Spark AQE (Xue et al., 2024) and QuerySplit (Zhao et al., 2023) adjust execution without leveraging learning-based guidance. In contrast, LQRS represents the first learned query optimizer on Spark SQL that jointly supports pre-execution and in-execution plan refinement, effectively unifying learned query optimization with adaptive query processing.
LQRS continuously collects runtime execution statistics and exposes them as plan features, including actual cardinality observed during query execution. These runtime features allow the optimizer to analyze partially executed plans using information unavailable before execution, enabling more effective optimization decisions during query processing. To enhance learning stability, LQRS adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner. In addition, LQRS unifies pre-execution and in-execution plan optimization within a single learning process, allowing in-execution learned knowledge to be leveraged for pre-execution optimization.
Figure 1. LQRS is the first learned query optimizer achieving both pre-execution and in-execution query plan optimization on Spark SQL.
[…]
Our contributions can be summarized as follows:
  • We design and implement LQRS, a learned query re-optimizer which incorporates reinforcement learning into the Spark AQE framework to make decisions during query execution based on actual runtime observations.
@@
  • We adopt a decision model built upon the TreeCNN architecture  (Mou et al., 2016), which learns optimization actions directly from query stage-level feedback collected during query execution. By using true cardinalities and size in bytes as input features, the model achieves better query performance with a simplified architecture.
@@
  • We design a plug-and-play planner extension built upon the extensibility interfaces of Spark SQL, enabling online plan modification during execution. The extension intercepts runtime execution plans, interacts with the decision model, and applies optimization actions such as join reordering.
@@
  • We conduct extensive experiments on four public benchmarks. Experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.
2. Related Work
Figure 2. This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
Learned Query Optimization
Since LEO (Stillger et al., 2001) in 2001 introduced the concept of a learned query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers. Plan-constructor methods aim to build query plans directly from scratch. These approaches enable the discovery of novel and potentially more efficient plans.
@@ 2. Related Work
Neo  (Marcus et al., 2019) pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle. Balsa  (Yang et al., 2022) further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques. LOGER  (Chen et al., 2023a) enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly. GLO  (Chen et al., 2024a) improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters.
@@
Plan-steerer methods improve query plans by exploiting the internal mechanisms of the native optimizer. Instead of constructing plans from scratch, they guide the optimizer to explore multiple candidate plans and select the one with the lowest estimated cost. For example, Bao  (Marcus et al., 2021), DeepO  (Sun et al., 2022), SCOPE  (Negi et al., 2021), FASTgres  (Woltmann et al., 2023), and AutoSteer  (Anneser et al., 2023) generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers. HybridQO  (Yu et al., 2022) explores alternative join orders by injecting leading hints. Lero (Chen et al., 2024b) modifies cardinality estimates in the native optimizer to produce diverse plans.
LQRS differs from existing LQOs in that it leverages runtime plan feature representations to make more informed optimization decisions and allows optimization decisions made pre-execution to be corrected during query execution.
@@
Adaptive query processing (AQP) has been extensively investigated to facilitate query plan adjustments during execution based on runtime feedback  (Amsaleg et al., 1996; Borovica-Gajic et al., 2015; Deshpande and Hellerstein, 2004; Deshpande et al., 2007; Kabra and DeWitt, 1998b; Eurviriyanukul et al., 2010; Zhu et al., 2017; Justen et al., 2024; Lyu et al., 2024). The primary AQP techniques are categorized as operator-based AQP and plan-based AQP.
Operator-based AQP adapts the behavior of physical operators during execution. For instance, Lookahead Information Passing (Zhu et al., 2017) generalizes semi-join techniques to improve the pipelining of equijoins. The Adaptive Join Algorithm (Ward, 2019) dynamically selects between hash join and nested loop join during runtime.
Plan-based AQP (or query re-optimization) re-invokes the optimizer at certain points during query execution to revise structural decisions such as join orders or the execution order of subqueries  (Kabra and DeWitt, 1998a; Markl et al., 2004; Neumann and Galindo-Legaria, 2013; Karanasos et al., 2014; Herlihy et al., 2024; Wang et al., 2023; Zhao et al., 2023). Following the pioneering work on dynamic re-optimization (Kabra and DeWitt, 1998a), Adaptive Metaprogramming (Herlihy et al., 2024) moves recursive query optimization and code generation from compile time to runtime, enabling customized join-order optimization based on runtime information. LPCE (Wang et al., 2023) adopts a two-stage learning strategy, consisting of an initial model and a refinement model, to progressively improve cardinality estimation accuracy for query operators. QuerySplit (Zhao et al., 2023) proposes a proactive query re-optimization framework that decomposes a query into multiple logical subqueries and determines their execution order according to estimated costs and output cardinalities.
While most AQP techniques are not learning-based, LQRS extends the plan-based AQP paradigm and leverages reinforcement learning to guide optimization decisions based on runtime feedback. It also integrates pre-execution and in-execution optimization into a unified learning process and supports adaptation at both the operator and plan levels.
3. Illustrative Overview
3.1. Spark SQL’s Query Execution Model
In Spark SQL, query execution is organized around query stages, which serve as the fundamental units of physical execution and runtime optimization. A query stage corresponds to a contiguous subgraph of the physical plan that can be executed without shuffle boundaries; stages are separated by shuffle operations that materialize intermediate results (for example, when a join is performed on a different join key from the preceding join, a shuffle is required to repartition the data, thereby introducing a new query stage). Once a query stage completes, Spark obtains accurate runtime statistics for that stage, including output cardinalities and data sizes, which are unavailable prior to execution.
3.2. Spark SQL’s Adaptive Query Execution
In order to tackle the challenges of cardinality estimation and physical plan cost estimation, Spark SQL 3.0 (2020) introduced AQE. AQE starts with an initial pre-execution query plan based on rough estimations and then re-optimizes it during execution using actual runtime observations.
Figure 2 illustrates a mechanism adopted by AQE called dynamic join selection. In Figure 2, the query is executed in multiple query stages. The red segments in the figure correspond to standard query execution without adaptation, while the blue and green parts indicate the two plan modifications introduced by AQE interventions at runtime. After the first query stage, which executes the join $t1\bowtie t2$ and produces only 5 rows and 70 bytes of intermediate result, AQE decides to switch the join strategy for the second stage from SMJ to BHJ. After the second stage, which executes $(t1\bowtie t2)\bowtie t3$ and produces 100,000 rows and 700,000 bytes of intermediate result, AQE again adapts the plan, switching the join strategy for the third stage from BHJ back to SMJ. By reasoning at the granularity of query stages during execution, AQE can make more informed decisions based on actual intermediate cardinalities rather than relying solely on rough pre-execution estimates.
3.3. Spark AQE’s Limitations
Despite its ability to adapt join strategies at runtime, Spark AQE has inherent limitations. First, the effectiveness of AQE’s dynamic join selection is constrained by human-configured thresholds. In particular, AQE only converts an SMJ to a BHJ if the size of one side’s intermediate result is smaller than the broadcast join threshold (BJT). Since broadcasting a large table can lead to worker crashes, administrators typically configure BJT with an over-conservative static value, which limits the performance benefits.
Second, AQE cannot change the join order once query execution begins. The optimizer relies on the pre-execution plan to determine the sequence of joins. As a result, queries with suboptimal join orders may still incur significant performance penalties, leaving untapped optimization potential that current AQE cannot address.